# Generalized Short Path Algorithms: Towards Super-Quadratic Speedup over Markov Chain Search for Combinatorial Optimization

### v1

@@ 1.3 Related Works
As discussed at length in the above, the primary inspiration for our work comes from the earlier short path algorithms developed in the series of works [Has18a, Has19b, Has18b, DPCB23]. We now discuss other related quantum algorithms and their connections to our results. The most closely related family of algorithms are based on discrete time quantum walks. The framework of Magniez at al [MNRS11] considers a framework that closely matches the one considered here. In [MNRS11], the authors seek to accelerate a classical algorithm, that first prepares in time $\mathrm{S}$ a sample from the stationary distribution of a Markov Chain with spectral gap $\delta$, that has $\epsilon$ overlap with some marked state that can be checked in time $\mathrm{C}$, and then runs the Markov Chain to repeatedly draw and check samples from the stationary distribution. The classical algorithm finds a marked element in time $\mathcal{O}\left(\mathrm{S}+\epsilon^{-1}(\delta^{-1}\mathrm{U}+\mathrm{C})\right)$ where $\mathrm{U}$ is the cost of simulating one step of the chain. The quantum algorithm obtains a runtime of $\mathcal{O}\left(\mathrm{S}+\epsilon^{-1/2}(\delta^{-1/2}\mathrm{U}+\mathrm{C})\right)$.
@@
In our setting $\epsilon$ corresponds to $\pi(E^{\ast})$ and falls exponentially, whereas $\mathrm{S,U,C},\delta^{-1}$ all grow polynomially. Applying the [MNRS11] framework, we obtain a quantum runtime of $\mathcal{O}^{\ast}(\pi(E^{\ast})^{0.5})$. If the conditions for our framework are met, we obtain a runtime of $\mathcal{O}^{\ast}(\pi(E^{\ast})^{0.5-c(n)})$. In this setting, our algorithm accelerates the [MNRS11] framework in a manner analogous to how [DPCB23] accelerates Grover search. Another common framework for analyzing search via quantum walk is that of Szegedy [Sze04] where the quadratic speedup is over the hitting time of a marked vertex. However, since the Hitting Time $\mathrm{HT}$ from stationarity satisfies $\epsilon^{-1}\leq\mathrm{HT}\leq\epsilon^{-1}\delta^{-1}$, we have $\mathrm{HT}=\mathcal{O}^{\ast}(\epsilon^{-1})$ in our setting where $\epsilon^{-1}$ is exponentially larger than $\delta^{-1}$, and the runtime of both quantum walk frameworks match upto polynomial factors.
@@
Aside from the works we have already mentioned, quantum algorithms have also been successfully leveraged to obtain speedups for solving linear systems of equations [HHL09, CKS17, CGJ19], computing partition functions [HW20, CH23], estimating the volume of convex bodies [CCH+23], as well as sampling from both log-concave [CLL+22] and non-log-concave [OLMW24] distributions. We note that our work is primarily concerned with exponential time search algorithms instead of the randomized approximation schemes considered in these papers. It would be interesting to understand whether our methods can be used to obtain super-quadratic speedups for exponential time counting algorithms [GLR21]. Along the line of super-quadratic quantum speedups, there is also recent work providing positive results for Tensor Principle Component Analysis [SOKB24], and approximation algorithms for combinatorial optimization [JSW+24].
@@ 2 Preliminaries
We also define $\widetilde{\mathcal{O}}(f(x))=\mathcal{O}(f(x)\cdot\textup{polylog}(f(x)))$ and $\mathcal{O}^{*}(2^{f(x)})=\mathcal{O}(2^{f(x)}\cdot\operatorname{poly}(x))$.
@@ 2.1 Probability toolbox
In what follows let $\mathcal{X}$ be a finite set satisfying $|\mathcal{X}|=V$. A Markov chain $\mathcal{M}$ is a random process that defines movements between elements of $\mathcal{X}$. Transitions between states are determined according to a fixed probability distribution, and can be represented by an $V\times V$ (though not necessarily symmetric) transition matrix $P$. The entry $P_{kj}:=P(k,j)$ is the probability of making a transition from $k$ to $j$, and the rows of $P$ sum to 1 to preserve normalization $\sum_{j\in\mathcal{X}}P(k,j)=1$; we say that such a matrix is stochastic.
@@
For a function $f:\mathcal{X}\mapsto\mathbb{R}$, we define
[…]
@@
$\textup{TV}(\mu,\nu)=\textup{sup}_{f:\lVert f\rVert\leq 1}\frac{1}{2}(\mathbb{E}_{\mu}[f(x)]-\mathbb{E}_{\nu}[f(x)]).$
@@
$11\geq\lambda_{1}\geq\cdots\geq\lambda_{N}\geq-$ 1 .
It is known that $\lambda_{1}=1$ and $\lambda_{2}<1$. We define the spectral gap of a reversible Markov chain to be:
@@
Let $f:\mathcal{X}\mapsto\mathbb{R}$. The Dirichlet form, $\mathcal{D}(f,f)$, generated by a Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$ is defined by
@@
and if $P$ is reverisble, then $I-P$ is symmetric with respect to the $\pi$-inner product, and so $\mathcal{D}$ extends to an inner product for functions $f,g$:
[…]
@@
A Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$ satisfies a log-Sobolev inequality with constant $\omega:=\omega_{LS}$ if
@@
A Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$ satisfies a log-Sobolev inequality with constant $\omega_{\textup{MLS}}$ if
$\mathcal{D}(f,\ln f)\geq\omega_{\text{MLS}}\textup{Ent}_{\pi}(f).$
@@
The Poincaré and log-Sobolev inequalities belong to a group known as the functional inequalities. They are well-defined for non-reversible chains, although Poincaré now bounds the singular-value gap, and enable one to bound the corresponding mixing time [Cha23].
@@
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain. The $P$-pseudo Lipschitz norm of $f:\mathcal{X}\mapsto\mathbb{R}$ is defined to be
[…]
@@ 2.2 Short path algorithms
Let $H:\left\{-1,1\right\}^{n}\mapsto\mathbb{R}$ be a classical cost function satisfying $\sum_{z}H(z)=0$ for $H$ with no constant term. Consider the combinatorial optimization problem
@@
Recently, Hastings [Has18a] and Dalzell et al. [DPCB23] proposed a new paradigm for avoiding the first-order phase transition problem with the QAA. Following the approach of [DPCB23], prototypical adiabatic optimization is eschewed through two modifications. First, the term $H$ is replaced with $g_{\eta}\left(\frac{H}{|E^{\ast}|}\right)$ for a piecewise-linear function $g_{\eta}:[-1,\infty)\mapsto[-1,0]$ parameterized by $\eta\in[0,1)$:
@@
leading to the Hamiltonian:
@@
where $X$ has been normalized by its spectral norm, and $b\in[0,\infty)$. Second, rather than slowly evolve from $-\frac{X}{n}$ to $\frac{H}{|E^{\ast}|}$ as in the QAA, we jump from $-\frac{X}{n}$ to the ground state |$\ket{\psi_{b}}$ of $H_{b}$ for some value of $b$ that is independent of $n$ (where the spectral gap is guaranteed to be large), and then jump from $H_{b}$ to the ground-state space of $\frac{H}{|E^{\ast}|}$. Note in [DPCB23] they also allow for scaling $H$ by an overestimate of $\lvert E^{*}\rvert$, for simplicity we just stick with scaling by $\lvert E^{*}\rvert$.
The jumps are accomplished using a unitary $U$, which combines phase estimation and amplitude amplification. For a high-level understanding, suppose we seek to enact a jump between two Hamiltonians $H_{1}$ and $H_{2}$, each acting on $n$ qubits. Denote the ground state of $H_{1}$ by |$\ket{\psi_{1}}$, and let $\Pi_{2}$ be the projector onto the ground space of $H_{2}$. The unitary $U$ first employs phase estimation to implement reflection operators $R_{1}$ and $R_{2}$ that reflect about the state |$\ket{\psi_{1}}$, and the groundspace of $H_{2}$, respectively. If $\delta_{j}$ is the spectral gap of Hamiltonian $H_{j}$, the operator $R_{j}$ can be implemented up to error $\varepsilon$ using $\delta_{j}^{-1}\log(1/\varepsilon)$ calls to a block-encoding of $H_{j}$, and often realizable using $\operatorname{poly}(n)$ gates. When $H_{j}$ is a classical Hamiltonian, $R_{j}$ can be implemented using $\operatorname{poly}(n)$ gates irrespective of $\delta_{j}$. From here, the unitary $U$ employs fixed-point amplitude amplification [YLC14] to implement $\frac{\Pi_{2}\ket{\psi_{1}}}{\left\|\Pi_{2}\ket{\psi_{1}}\right\|}$, requiring at most $\mathcal{O}\left(\left\|\Pi_{2}\ket{\psi_{1}}\right\|^{-1}\log(1/\varepsilon)\right)$ applications each of $R_{1}$ and $R_{2}$.
The algorithm is initialized to |+:=|$\ket{\boldsymbol{+}}:=\ket{+}^{\otimes n}$, the ground state of $-\frac{X}{n}$. Then, the ground state |$\ket{\psi_{b}}$ of $H_{b}$ is prepared by jumping from $-X/n$ to $H_{b}$. Finally, we prepare $\frac{\Pi^{\ast}\ket{\psi_{b}}}{\lVert\Pi^{\ast}\ket{\psi_{b}}\rVert}$ by jumping from $H_{b}$ to the classical Hamiltonian $\frac{H}{{\left|E^{\ast}\right|}}$. The state $\Pi^{\ast}\ket{\psi_{b}}$ is a superposition of optimal solutions to $\min_{z\in\{\pm 1\}^{n}}H(z)$, and thus measurement in the computational basis yields an optimal bit-string $z^{\ast}$ with high probability. The first jump is small (in the sense that the success probability is nearly 1), whereas the second jump is large (the success probability is exponentially small). The resulting time complexity scales as $\mathcal{O}^{*}\left(2^{\left(\frac{1}{2}-c\right)n}\right)$, indicating a super-Grover speedup when $c>0$. In [Has18a], the order of short and long jump steps is reversed. We refer to both approaches [Has18a, DPCB23] as short path algorithms.
@@ 3.1 Framework
Define $\mathcal{X}\subseteq\{-1,1\}^{n}$ and let $H\colon\mathcal{X}\mapsto\mathbb{R}$ be a cost function. We are interested in (exactly) determining $z^{\ast}\in\mathcal{X}$ such that
@@
We also use $H$ to refer to a diagonal Hamiltonian in a Hilbert space with a basis indexed by $z\in\mathcal{X}$, with $\langle z|H|z\rangle$ for the Hamiltonian identified with $H(z)$. Whether we are referring to the quantum Hamiltonian or the function will be clear from context. We define $E^{\ast}:=\min_{z\in\mathcal{X}}H(z)$ and assume everywhere that the cost is scaled to ensure $E^{\ast}<0$. We will further assume that $|\mathcal{X}|$ is super-polynomial, as our primary concern is with super-quadratic speedups over exponential time algorithms. If $\pi$ is a distribution such that $\pi(E^{\ast})$ is the probability that a sample from $\pi$ is a global minimizer and there exists a Markov Chain with transition matrix $P$ that mixes to stationary distribution $\pi$ such the mixing time is bounded by $t_{\mathrm{mix}}(\varepsilon)=\operatorname{poly}(n,\log(\varepsilon^{-1}))$.
@@
Prerequisites: Solution space $\mathcal{X}\subset\{-1,1\}^{n}$, Cost function $H\colon\mathcal{X}\mapsto\mathbb{R}$, distribution $\pi$ such that $\pi(E^{\ast})$ is the probability that a sample from $\pi$ is a global minimizer, a Markov Chain with transition matrix $P$ and mixing time $t_{\mathrm{mix}}(\pi(E^{\ast})/2)=\mathcal{O}(\operatorname{poly}(n))$.
      Input: Description of $\operatorname{poly}(n)$ time procedures to evaluate $H$ and perform a step of the Markov Chain described by $P$, failure probability $\delta$.
@@
2:while $i\leq\frac{2}{\pi(E^{\ast})}\log\left(\frac{1}{\delta}\right)$ do
3:     Simulate $t_{\mathrm{mix}}(n)$ steps of $P$ to obtain sample $\tilde{z}$
@@
Under these conditions, it follows that Algorithm 1 finds the global minimizer of $H$ in time $\mathcal{O}^{\ast}\left((\pi(E^{\ast}))^{-1}\right)$. Our framework seeks to accelerate this runtime, and so we assume the same setting for the quantum algorithm. In order to define the quantum framework we must access $H,P,$ and $\pi$. We assume the existence of the following subroutines:
@@
The block-encoding for $H_{b}$ can be constructed using the linear-combination-of-block-encodings technique [GSLW19] using $U_{D(P)}$ and $U_{H}$.
The framework is specified in Algorithm 2. The implementation of the jumps uses the framework from [DPCB23, Proposition 21] that performs fixed point amplitude amplification using reflections constructed from the block-encodings of the Hamiltonians and the Quantum Singular Value Transform. The overall runtime of Algorithm 2, in terms of the number of queries to $U_{\pi}$ and block-encodings of $H$ and $D(P)$, is
@@
1:Prepare |$\ket{\sqrt{\pi}}$, the ground state of $-D(P)$.
2:Short Jump: Prepare |$\ket{\psi_{b}}$ up to exponentially small error with jump $-D(P)\rightarrow H_{b}$.
3:Long Jump: Prepare $\frac{\Pi^{\ast}\ket{\psi_{b}}}{\lVert\Pi^{\ast}\ket{\psi_{b}}\rVert}$ up to exponentially small error with jump $H_{b}\rightarrow\frac{H}{{\left|E^{\ast}\right|}}$.
Our bounds on the runtime rely on two conditions, that we view as uplifted versions of corresponding notions in [DPCB23] to the case of general Markov Chains. Our first conditons captures the smoothness of the cost function under applications of the transition matrix.
@@
where $h_{\eta}(x):=g_{\eta}(\frac{x}{\lvert E^{*}\rvert})$. If the conditions holds for all $10<\eta<$ 1 we omit it and simply say $H$ is $\Delta_{P}$ stable under $\mathcal{M}$.
The analysis in the following sections will clarify that $\Delta_{P}$-stable is a generalization of the $\alpha$-subdepolarizing condition introduced in [DPCB23]. In fact, it is equivalent to a more syntactically obvious generalization of the $\alpha$-subdepolarizing condition, we state Definition 3.1 as the primary condition as it is easier to demonstrate and interpret in most cases.
@@
$\pi(E\leq(1-\eta)E^{\ast})\leq\pi(E^{\ast})^{\gamma}.$
We are now ready to state the main results concerning our framework. These results are subject to further technical conditions on the Markov Chain used for the search, and are formulated in terms of conditions that lead to efficient mixing time bounds. We have two variants of our result, the first relies on a logarithmic-Sobolev inequality for $P$. We have the following result:
Theorem 3.3 (informal).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible Markov chain, and let $H:\mathcal{X}\mapsto\mathbb{R}$ be a diagonal, $\Delta_{P}$-stable Hamiltonian with ground state energy $E^{\ast}$, that satisfies the $\gamma$ spectral density condition for some parameters $\eta,\gamma$. In addition, suppose $\mathcal{M}$ satisfies an $\omega$ log-Sobolev inequality. If $\omega^{-1}=\Theta(\ln(1/\pi(E^{\ast}))$, then there exists a constant $b$, such that under Assumption 1, Algorithm 2 determines the ground state of $H$ over $\mathcal{X}$ with running time
We also present a variant of the above result that only relies on the weaker Poincaré inequality.
Theorem 3.4 (informal).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversbile Markov chain, and let $H:\mathcal{X}\mapsto\mathbb{R}$ be a diagonal, $\Delta_{P}$-stable Hamiltonian with ground state energy $E^{\ast}$, that satisfies a spectral density condition. In addition, suppose $\mathcal{M}$ satisfies a $\delta$ Poincaré inequality. If $\delta$ is independent of the problem size, then there exists a constant $b$, such that under Assumption 1, Algorithm 2 determines the ground state of $H$ over $\mathcal{X}$ with running time
The spectral density condition with respect to non-uniform starting states presents a technical challenge, since unlike the uniform distribution over all bitstrings of length $n$, they are no longer product measures. Fortunately, the condition that $\pi$ is the ground state of a fast mixing Markov Chain allows for simplification via concentration inequalities for Markov chains, e.g. the so-called Herbst argument [Lal13], that allows the spectral density conditions to be established as long as the cost function satisfies has an appropriately bounded pseudo-Lipschitz norm $\lVert H\rVert_{P}$. Theorems 3.3 and 3.4 are proven in Section 4, in order to establish specializations that rely on a bounded pseudo-Lipschitz norm.

### v2

@@ 1.3 Related Works
As discussed at length in the paragraph above, the primary inspiration for our work comes from the earlier short path algorithms developed in the series of works [Has18a, Has19, Has18b, DPCB23]. We now discuss other related quantum algorithms and their connections to our results. The most closely related family of algorithms are based on discrete time quantum walks. The framework of Magniez at al [MNRS11] considers a framework that closely matches the one considered here. In [MNRS11], the authors seek to accelerate a classical algorithm, that first prepares in time $\mathrm{S}$ a sample from the stationary distribution of a Markov chain with spectral gap $\delta$, that has $\epsilon$ overlap with some marked state that can be checked in time $\mathrm{C}$, and then runs the Markov chain to repeatedly draw and check samples from the stationary distribution. The classical algorithm finds a marked element in time $\mathcal{O}\left(\mathrm{S}+\epsilon^{-1}(\delta^{-1}\mathrm{U}+\mathrm{C})\right)$ where $\mathrm{U}$ is the cost of simulating one step of the chain. The quantum algorithm obtains a runtime of $\mathcal{O}\left(\mathrm{S}+\epsilon^{-1/2}(\delta^{-1/2}\mathrm{U}+\mathrm{C})\right)$.
@@
In our setting $\epsilon$ corresponds to $\pi(E^{\ast})$ and falls exponentially, whereas $\mathrm{S,U,C},\delta^{-1}$ all grow polynomially. Applying the [MNRS11] framework, we obtain a quantum runtime of $\mathcal{O}^{\ast}(\pi(E^{\ast})^{-0.5})$. If the conditions for our framework are met, we obtain a runtime of $\mathcal{O}^{\ast}(\pi(E^{\ast})^{-0.5+c(n)})$. In this setting, our algorithm accelerates the [MNRS11] framework in a manner analogous to how [DPCB23] accelerates Grover search. Another common framework for analyzing search via quantum walk is that of Szegedy [Sze04] where the quadratic speedup is over the hitting time of a marked vertex. However, since the Hitting Time $\mathrm{HT}$ from stationarity satisfies $\epsilon^{-1}\leq\mathrm{HT}\leq\epsilon^{-1}\delta^{-1}$, we have $\mathrm{HT}=\mathcal{O}^{\ast}(\epsilon^{-1})$ in our setting where $\epsilon^{-1}$ is exponentially larger than $\delta^{-1}$, and the runtime of both quantum walk frameworks match up to polynomial factors.
@@
Aside from the works we have already mentioned, quantum algorithms have also been successfully leveraged to obtain speedups for solving linear systems of equations [HHL09, CKS17, CGJ19], computing partition functions [HW20, CH23], estimating the volume of convex bodies [CCH+23], as well as sampling from both log-concave [CLL+22] and non-logconcave [OLMW24] distributions. We note that our work is primarily concerned with exponential time search algorithms instead of the randomized approximation schemes considered in these papers. It would be interesting to understand whether our methods can be used to obtain super-quadratic speedups for exponential time counting algorithms [GLR21]. Along the line of super-quadratic quantum speedups, there is also recent work providing positive results for Tensor Principle Component Analysis [SOKB24], and approximation algorithms for combinatorial optimization [JSW+24].
@@ 2 Preliminaries
We will use $\mathcal{G}_{\text{RG}}(n,d)$ to denote the uniform random ensemble of regular graphs on $n$ vertices with degree $d$. Also, $\mathcal{G}_{\text{ER}}(n,p)$ will be used to denote the Erdős-Rényi ensemble, where there are $n$ vertices and an edge is created between two vertices with probability $p$.
@@
We also define $\widetilde{\mathcal{O}}(f(x))=\mathcal{O}(f(x)\cdot\textup{polylog}(x))$ and $\mathcal{O}^{*}(2^{f(x)})=\mathcal{O}(2^{f(x)}\cdot\operatorname{poly}(x))$.
@@ 2.1 Probability Toolbox
In what follows let $\mathcal{X}$ be a finite set satisfying $|\mathcal{X}|=V$. A Markov chain $\mathcal{M}$ is a random process that defines movements between elements of $\mathcal{X}$. Transitions between states are determined according to a fixed probability distribution, and can be represented by an $V\times V$ (though not necessarily symmetric) transition matrix $P$. The entry $P_{kj}:=P(k,j)$ is the probability of making a transition from $k$ to $j$, and the rows of $P$ sum to 1 to preserve normalization $\sum_{j\in\mathcal{X}}P(k,j)=1$; we say that such a matrix is stochastic. Additionally, $\mathcal{M}$ is ergodic if it is aperiodic (it does not get trapped in cycles) and irreducible (every state can reach every other).
@@
For a function $f:\mathcal{X}\rightarrow\mathbb{R}$, we define
$\mathbb{E}_{y\underset{P}{\sim}x}[f(x)]:=(Pf)(x)=\sum_{y\in\mathcal{X}}P(x,y)f(y),$ (1)
@@
$\textup{TV}(\mu,\nu)=\textup{sup}_{f:\lVert f\rVert_{\infty}\leq 1}\frac{1}{2}(\mathbb{E}_{\mu}[f(x)]-\mathbb{E}_{\nu}[f(x)]).$
@@
$1\geq\lambda_{1}\geq\cdots\geq\lambda_{N}\geq-1.$
It is known that $\lambda_{1}=1$ and $\lambda_{2}<1$. If the chain is aperiodic, then we can take $\lambda_{N}>-1$, where any chain can be made aperiodic by introducing self-loops, i.e. making the chain lazy. We define the spectral gap of a reversible Markov chain to be:
@@
Let $f:\mathcal{X}\rightarrow\mathbb{R}$. The Dirichlet form, $\mathcal{D}(f,f)$, generated by a Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$ is defined by
@@
and if $P$ is reversible, then IPI-P is symmetric with respect to the $\pi$-inner product, and so $\mathcal{D}$ extends to an inner product for functions $f,g$:
$\mathcal{D}(f,g)=\langle f,(I-P)g\rangle_{\pi}=\frac{1}{2}\mathbb{E}_{x\sim\pi}\mathbb{E}_{y\underset{P}{\sim}x}[(f(x)-f(y))(g(x)-g(y))].$
@@
A Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$ satisfies a log-Sobolev inequality with constant $\omega:=\omega_{\textup{LS}}$ if
@@
A Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$ satisfies a modified log-Sobolev inequality with constant $\omega_{\textup{MLS}}$ if
$\mathcal{D}(f,\ln f)\geq\omega_{\textup{MLS}}\textup{Ent}_{\pi}(f).$
@@
The Poincaré and log-Sobolev inequalities belong to a group known as the functional inequalities. They are also well-defined for non-reversible chains, although Poincaré now bounds the singular-value gap, and enable one to bound the corresponding mixing time [Cha23].
@@
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain. The $P$-pseudo Lipschitz norm of $f:\mathcal{X}\rightarrow\mathbb{R}$ is defined to be
$\lVert f\rVert_{P}:=\max_{x\in\mathcal{X}}\mathbb{E}_{y\underset{P}{\sim}x}[(f(x)-f(y))^{2}].$
@@ 2.2 Short Path Algorithms
Let $H:\left\{-1,1\right\}^{n}\rightarrow\mathbb{R}$ be a classical cost function satisfying $\sum_{z}H(z)=0$ for $H$ with no constant term. Consider the combinatorial optimization problem
@@
Recently, Hastings [Has18a] and Dalzell et al. [DPCB23] proposed a new paradigm for avoiding the first-order phase transition problem with the QAA. Following the approach of [DPCB23], prototypical adiabatic optimization is eschewed through two modifications. First, the term $H$ is replaced with $g_{\eta}\left(\frac{H}{|E^{\ast}|}\right)$ for a piecewise-linear function $g_{\eta}:[-1,\infty)\rightarrow[-1,0]$ parameterized by $\eta\in[0,1)$:
@@
leading to the short-path Hamiltonian:
@@
[…]
@@ 3.1 Framework
Define $\mathcal{X}\subseteq\{-1,1\}^{n}$ and let $H\colon\mathcal{X}\rightarrow\mathbb{R}$ be a cost function. We are interested in (exactly) determining $z^{\ast}\in\mathcal{X}$ such that
@@
We also use $H$ to refer to a diagonal Hamiltonian in a Hilbert space with a basis indexed by $z\in\mathcal{X}$, with $\langle z|H|z\rangle$ for the Hamiltonian identified with $H(z)$. Whether we are referring to the quantum Hamiltonian or the function will be clear from context. We define $E^{\ast}:=\min_{z\in\mathcal{X}}H(z)$ and assume everywhere that the cost is scaled to ensure $E^{\ast}<0$. We will further assume that $|\mathcal{X}|$ is super-polynomial, as our primary concern is with super-quadratic speedups over algorithms with at least super-polynomial runtime. Suppose $\pi$ is a distribution such that $\pi(E^{\ast})$ is the probability that a sample from $\pi$ is a global minimizer. Suppose there exists a Markov chain with transition matrix $P$ that mixes to stationary distribution $\pi$, then the mixing time is bounded by $t_{\mathrm{mix}}(\varepsilon)=\operatorname{poly}(n,\log(\varepsilon^{-1}))$.
@@
Prerequisites: Solution space $\mathcal{X}\subset\{-1,1\}^{n}$, Cost function $H\colon\mathcal{X}\rightarrow\mathbb{R}$, distribution $\pi$ such that $\pi(E^{\ast})$ is the probability that a sample from $\pi$ is a global minimizer, a Markov chain with transition matrix $P$ and mixing time $t_{\mathrm{mix}}=t_{\mathrm{mix}}(\pi(E^{\ast})/2)=\mathcal{O}(\operatorname{poly}(n))$.
   Input: Description of $\operatorname{poly}(n)$ time procedures to evaluate $H$ and perform a step of the Markov chain described by $P$, failure probability $\epsilon$.
@@
2:while $i\leq\frac{2}{\pi(E^{\ast})}\log\left(\frac{1}{\epsilon}\right)$ do
3:  Simulate $t_{\mathrm{mix}}$ steps of $P$ to obtain sample $\tilde{z}$
@@
Under these conditions, it follows that Algorithm 1 finds the global minimizer of $H$ in time $\mathcal{O}^{\ast}\left(\pi(E^{\ast})^{-1}\right)$. Our framework seeks to accelerate this runtime, and so we assume the same setting for the quantum algorithm. In order to define the quantum framework we must access $H,P,$ and $\pi$. We assume the existence of the following subroutines:
@@
The block-encoding for $H_{b}$ can be constructed using the linear combination of block-encodings technique [GSLW19] using $U_{D(P)}$ and $U_{H}$.
The framework is specified in Algorithm 2. The implementation of the jumps uses the framework from [DPCB23, Proposition 21] that performs fixed point amplitude amplification using reflections constructed from the block-encodings of the Hamiltonians and the quantum singular value transformation (QSVT). The overall runtime of Algorithm 2, in terms of the number of queries to $U_{\pi}$ and block-encodings of $H$ and $D(P)$, is
@@
[…]
Our bounds on the runtime rely on two conditions, that we view as uplifted versions of corresponding notions in [DPCB23] to the case of general Markov chains. Our first conditon captures the smoothness of the cost function under applications of the transition matrix.
@@
$\mathbb{E}_{y\underset{P}{\sim}x}[h_{\eta}\left(H(y)\right)]\leq h_{\eta}\left(H(x)+\Delta_{P}(\eta)\right),\quad\forall x\in\mathcal{X}$ (10)
@@
The analysis in the following sections will clarify that $\Delta_{P}$-stability is a generalization of the $\alpha$-subdepolarizing condition introduced in [DPCB23]. In fact, it is equivalent to a more syntactically obvious generalization of the $\alpha$-subdepolarizing condition, we state Definition 3.1 as the primary condition as it is easier to demonstrate and interpret in most cases.
@@
$\pi(E\leq(1-\eta)E^{\ast})\leq\pi(E^{\ast})^{\gamma},$
for some $\eta\in(0,1)$.
In section 4.2, we will present conditions on the underlying Markov chain that enable bounding $\gamma$. However, in some cases, we will simply assume $\gamma=\Theta(1)$, which as mentioned earlier is a mild condition for hard problems.
We are now ready to state the main results concerning our framework. These results are subject to further technical conditions on the Markov chain used for the search, and are formulated in terms of conditions that lead to efficient mixing time bounds. We have two variants of our result, the first relies on a logarithmic-Sobolev inequality for $P$.
Theorem 3.3 (Informal Theorem 4.2).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible, aperiodic Markov chain, and let $H:\mathcal{X}\rightarrow\mathbb{R}$ be a diagonal, $\Delta_{P}$-stable Hamiltonian with ground state energy $E^{\ast}$, that satisfies the $\gamma$ spectral density condition for some parameter $\eta$. In addition, suppose $\mathcal{M}$ satisfies an $\omega$ log-Sobolev inequality.Then, under Assumption 1, Algorithm 2 determines the ground state of $H$ over $\mathcal{X}$ with running time
$\mathcal{O}\left(\textup{poly}(n)\omega^{-1}[\pi(E^{\ast})^{-1}]^{\left(\frac{1}{2}-\frac{\eta(1-\eta)\lvert E^{\ast}\rvert b}{2\ln(1/\pi(E^{\ast}))\Delta_{P}}\right)}\right),$
where $b=\mathcal{O}\left(\gamma\omega\ln\left(\frac{1}{\pi(E^{\ast})}\right)\right)$.
The spectral density condition with respect to non-uniform starting states presents a technical challenge, since unlike the uniform distribution over all bitstrings of length $n$, they are no longer product measures. Fortunately, the condition that $\pi$ is the ground state of a fast mixing Markov chain allows for simplification via concentration inequalities for Markov chains, e.g. the so-called Herbst argument [Lal13], that allows the spectral density conditions to be established as long as the cost function has an appropriately bounded pseudo-Lipschitz norm $\lVert H\rVert_{P}$.
We also present a variant of the above result that only relies on the weaker Poincaré inequality and does not require the spectral density condition. Still, Corollary 4.9 shows that a Poincaré inequality does imply the spectral density condition, anyways, for some potentially falling $\gamma$.
Theorem 3.4 (informal Theorem 4.3).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible, aperiodic Markov chain, and let $H:\mathcal{X}\rightarrow\mathbb{R}$ be a diagonal, $\Delta_{P}$-stable Hamiltonian with ground state energy $E^{\ast}$ that satisfies the $\gamma$ spectral density condition for some parameter $\eta$. In addition, suppose $\mathcal{M}$ satisfies a $\delta$ Poincaré inequality. Then, under Assumption 1, Algorithm 2 determines the ground state of $H$ over $\mathcal{X}$ with running time
$\mathcal{O}\left(\textup{poly}(n)\delta^{-1}[\pi(E^{\ast})^{-1}]^{\left(\frac{1}{2}-\frac{\eta(1-\eta)\lvert E^{\ast}\rvert b}{2\ln(1/\pi(E^{\ast}))\Delta_{P}}\right)}\right),$
where $b=\mathcal{O}(\delta)$.
Theorems 3.3 and 3.4 are proven in Section 4, in order to establish specializations that rely on a bounded pseudo-Lipschitz norm. Additionally, in Section 4 we present conditions under which the super-quadratic speedup is asymptotic, i.e. does not fall with problem size $n$.